OpenStudy (anonymous):

The vertex of the graph of a quadratic function is (3, 1) and the y-intercept is (0, 10). WHat is the equation of the function?

OpenStudy (anonymous):

in vertex form it looks like \[f(x)=a(x-3)^2+1\] the only thing you don't know is \(a\) but since \((0,10)\) is on the graph you know \[f(0)=a(0-3)^2+1=10\] you can solve that for \(a\)

OpenStudy (anonymous):

how though i have no clue what im evan doin

OpenStudy (anonymous):

how do you solve for \(a\) ?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

k lets do it step by step what is \(0-3\) ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

right now what is \((-3)^2\) ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

k good so now \[a(0-3)^2+1=10\] is \[9a+1=10\] and we take two steps to solve for \(a\) 1) subtract 1 2) divide by 9

OpenStudy (anonymous):

ok so 1?

OpenStudy (anonymous):

yes \(a=1\) so \[f(x)=(x-3)^2+1\]

OpenStudy (anonymous):

on account of you don't really need to write \[f(x)=1(x-3)^2+1\]

OpenStudy (anonymous):

if you want, you could multiply out and combine like terms, or just leave it in "vertex form"

OpenStudy (anonymous):

thats not one of the answers tho?

OpenStudy (anonymous):

did you multiply out? \[(x-3)^2+1=(x-3)(x-3)+1\] etc

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

i dont remember...

OpenStudy (anonymous):

ok thanks though satellite73 it was really helpfull